Solve for the general solution using method of undetermined coefficients D4-1y=e-x

1
Expert's answer
2022-01-17T18:36:39-0500

Find solution of the associated homogeneous equation:


"(D^4-1)y=0"

Characteristic (auxiliary) equation is


"r^4-1=0"

"(r-1)(r+1)(r-i)(r+i)=0"

"r_1=-1, r_2=1, r_3=-i, r_4=i"

The general solution of the associated homogeneous equation is


"y_h=c_1e^x+c_2e^{-x}+c_3\cos x+c_4 \sin x"

Find the particular solution of the non homogeneous differential equation


"y_p=Axe^{-x}""y_p'=Ae^{-x}-Axe^{-x}"

"y_p''=-2Ae^{-x}+Axe^{-x}"

"y_p'''=3Ae^{-x}-Axe^{-x}"

"y_p''''=-4Ae^{-x}+Axe^{-x}"

Substitute


"-4Ae^{-x}+Axe^{-x}-Axe^{-x}=e^{-x}"

"A=-\dfrac{1}{4}"

The particular solution of the non homogeneous differential equation is


"y_p=-\dfrac{1}{4}xe^{-x}"

The general solution of the non homogeneous equation is


"y=c_1e^x+c_2e^{-x}+c_3\cos x+c_4 \sin x-\dfrac{1}{4}xe^{-x}"